### v2

@@ 3.1 Some preprocessing
It is known that every graph can be embedded into a 3-book [13, Proposition 3.1]. So we can without loss of generality assume that $\mathscr{C}$ contains no 3-book. We remove all the isolated vertices of $\mathscr{C}$, and remove the same number of isolated vertices of $G$ (to the extent possible); this does not affect whether $G$ embeds into $\mathscr{C}$. We then replace each isolated vertex of $G$ with an isolated edge; since $\mathscr{C}$ has no more isolated vertex, this does not affect embeddability of $G$ into $\mathscr{C}$. Finally, for the same reason, if $G$ contains at least two connected components that are paths, we replace all these components with a single edge. ∎
@@ 3.2 Structure of 2-complexes without 3-book or isolated vertex
Let $\mathscr{C}$ be a 2-complex without 3-book or isolated vertex, and let $p$ be a vertex of $\mathscr{C}$. Following [13, Section 2.2], we describe the possible neighborhoods of $p$ in $\mathscr{C}$. A cone at $p$ is a cyclic sequence of triangles $t_{1},\ldots,t_{k},t_{1}$ ($k\geq 3$), all incident to $p$, such that, for each $i=1,\ldots,k$, the triangles $t_{i}$ and $t_{i+1}$ (where $t_{k+1}=t_{1}$) share an edge incident with $p$, and any other pair of triangles have only $p$ in common. A corner at $p$ is a sequence of distinct triangles $t_{1},\ldots,t_{k}$, all incident to $p$, such that, for each $i=1,\ldots,k-1$, the triangles $t_{i}$ and $t_{i+1}$ share an edge incident with $p$, any other pair of these triangles have only $p$ in common, and no other triangle in $\mathscr{C}$ shares an edge incident with $p$ and belonging to one of $t_{1},\dots,t_{k}$. An isolated edge at $p$ is an edge incident to $p$ but not incident to any triangle. The cones, corners, and isolated edges at $p$ form the link components at $p$.
@@ 3.3 Topological data structure for 2-complexes
Given a 2-complex $\mathscr{C}$ without 3-book or isolated vertex, described by vertices, edges, and triangles and their incidence relations, we can easily compute a representation of $\mathscr{C}$ in that data structure, in polynomial time: Indeed, by ignoring the incidences created by vertices, we easily build a triangulation of the surface $\mathscr{S}$ (possibly disconnected, possibly with boundary) and a list of segments $S$; we then compute the topology of $\mathscr{S}$; finally, we mark the singular vertices, which are the vertices with several occurrences on $\mathscr{S}$ and/or on $S$. We remark that the size of the resulting data structure is at most linear in the number of vertices, edges, and triangles of the 2-complex $\mathscr{C}$, because, by Euler’s formula, any triangulated surface (possibly with boundary) with $k$ simplices has genus $O(k)$ and a number of boundary components that is $O(k)$. Thus, in the rest of this article, without loss of generality, we implicitly assume that $\mathscr{C}$ is given in the form of the above topological data structure. (Conversely, it is not hard to see that every 2-complex is homeomorphic to a 2-complex whose number of simplices that is linear in its size, but we will not need this fact.)
@@ 3.4 Proper and cellular graph embeddings on 2-complexes
Traditional data structures for graphs on surfaces handle graphs embedded cellularly; rotation systems [29] constitute one example of such a data structure. In order to have efficient algorithms, refined data structures, e.g., with graph-encoded maps [24] (see also [16, Section 2]), are needed. The basic element in the graph-encoded map is the flag, an incidence between a vertex, an edge, and a face of the graph. Three involutions allow to move from each flag to a nearby flag. Each flag contains a pointer to the underlying vertex, edge, and face.
One can easily extend such data structures to possibly non-cellular embeddings on surfaces [12, Section 2.2]. In this framework, one must store the topology of each face, which is not necessarily homeomorphic to a disk. Also, a face may have several boundary components; two-way pointers connect each face to one flag of each boundary component (or to an isolated vertex of the graph, if that boundary component is a single vertex); if a face is orientable and has several boundary components, then these pointers must induce a consistent orientation of these boundary cycles. It is important to remark that this data structure also allows to recover the topology of the underlying surface.
@@ 4 Partitioning graphs
  • Each face of $\Pi$ is labelled by an integer in $\{0,\ldots,k\}$ as follows: faces of $\Pi$ containing an edge in $E_{i}$ are labelled $i$, and the other ones are labelled 00. Since $\Pi$ contains the singular vertices of $\mathscr{C}$ and $\text{mid}(E_{1},\ldots,E_{k})$, and by the construction of the edges of $\Pi$, this labelling is well defined. We remark that this definition is valid also for faces of $\Pi$ that are parts of an isolated edge of $\mathscr{C}$.
@@
  • similarly, for each internal node $\nu$ of $\mathcal{B}$, the graph $\Pi(\Gamma,\nu)$ has at most $23(75c+27w)/$ 2 edges.
@@ 4.1 Monogons and bigons
A monogon of a graph $\Pi$ embedded on a 2-complex $\mathscr{C}$ is a face of $\Pi$ that is an open disk whose boundary is a single edge of $\Pi$ (a loop). Similarly, a bigon of $\Pi$ is a face of $\Pi$ that is an open disk whose boundary is the concatenation of two edges of $\Pi$ (possibly the same edge appearing twice). The following general lemma on graphs embedded on surfaces without monogons or bigons will be used below, and also in Section 8; some particular cases have been used before [8, Lemma 2.1].
@@
It is clear that the statement of the lemma holds whenever we are in one of the four above cases. So we can assume without loss of generality that each face of $\Pi^{\prime}$ is a triangle. Since $V(\Pi)=V(\Pi^{\prime})$ and $|E(\Pi)|\leq|E(\Pi^{\prime})|$, it suffices to prove the result for $\Pi^{\prime}$ instead of $\Pi$. Double-counting the incidences between edges and faces implies that the number of triangles $\tau$ satisfies $3\tau=2|E(\Pi^{\prime})|$; plugging this into Euler’s formula implies that $|V(\Pi^{\prime}|-|E(\Pi^{\prime})|/3=2-g$, so $|E(\Pi^{\prime})|=3g+3|V(\Pi^{\prime})|-6$, as desired. ∎
@@ 4.3 Rearranging $\Gamma$ with respect to an edge partition
Now, let us assume that $\Pi$ contains a sequence of bigons $B_{1},\ldots,B_{8}$ such that $B_{i}$ and $B_{i+1}$ share an edge for each $i$. So without loss of generality, we can assume that $B_{1}$ and $B_{5}$ are labelled $1$, $B_{3}$ and $B_{7}$ are labelled $2$, and the other bigons are labelled 00. We modify $\Gamma$ by exchanging the parts of $\Gamma$ inside $B_{3}$ and $B_{5}$; see Figure 4.5. The resulting embedding $\Gamma^{\prime}$ is also proper and cellular. This operation simplifies the endpoints $u$ and $v$ of these bigons. (If $u=v$, this is Case 2 from the definition of a simplification.)
@@
Finally, there remains to prove that, for each internal node $\nu$ of $\mathcal{B}$, the graph $\Pi(\Gamma,\nu)$ has at most $23(75c+27w)/$ 2 edges. Let $(E_{1},E_{2},E_{3})$ be the edge partition associated with $\nu$. By the claim we have just proved, it suffices to prove that the number of edges of this graph is at most half the sum of the numbers of edges of $\Pi(\Gamma,E_{1},E_{2}\cup E_{3})$, $\Pi(\Gamma,E_{1}\cup E_{2},E_{3})$, and $\Pi(\Gamma,E_{1}\cup E_{3},E_{2})$.
@@ 5.1 Bounding graphs
We compute our exhaustive list $\mathcal{L}_{\alpha}$ of sparse bounding graphs for $G_{\alpha}$ as follows. Initially, let this list be empty. Using Lemma 3.3, we enumerate all combinatorial maps $\Pi$ of graphs with at most $c+3w$ vertices and $23(75c+27w)/$ 2 edges properly embedded on $\mathscr{C}$ (possibly non-cellularly), with faces labelled 0, 1, 2, or 3, and such that the labels appearing on the vertices are exactly the vertices of the middle set of $\alpha$, $\beta$, or $\gamma$ (and each label appears exactly once). This takes $(c+w)^{O(c+w)}$ time. Whenever $\Pi_{1,3}\in\mathcal{L}_{\beta}$ and $(\Pi_{1,2})^{-}\in\mathcal{L}_{\gamma}$, we add $\Pi_{2,3}$ to $\mathcal{L}_{\alpha}$.
@@ 6 Reduction to proper cellular embeddings
The following result is not hard and essentially folklore (a related but slightly weaker result is provided by Matoušek, Sedgwick, Tancer, and Wagner [26, Lemma 3.1]), but we could not find a precise reference.
@@
    Case where $\gamma$ is non-separating but two-sided. This is only possible if $g\geq 2$. The cutting operation on $\mathscr{S}$ results in a single surface $\mathscr{S}^{\prime}$ with genus $g-2$. If $\mathscr{S}$ is orientable, then so is $\mathscr{S}^{\prime}$; otherwise, $\mathscr{S}^{\prime}$ is either orientable or non-orientable (unless of course $g=2$, in which case it is necessarily orientable, or $g$ is odd, in which case it is necessarily non-orientable). All surfaces $\mathscr{S}^{\prime}$ satisfying these constraints can be obtained.
@@
    Case where $\gamma$ is one-sided. This is only possible if $\mathscr{S}$ is non-orientable and $g\geq 1$. The cutting operation on $\mathscr{S}$ results in a single surface $\mathscr{S}^{\prime}$ with genus $g-1$, orientable or not (unless of course $g=1$, in which case it is orientable, or $g$ is even, in which case it is non-orientable). All surfaces $\mathscr{S}^{\prime}$ satisfying these constraints can be obtained.∎
@@
Let $\varphi(\mathscr{S}^{\prime})$ be equal to twice the total genus of $\mathscr{S}^{\prime}$ minus its number of connected components. By Lemma 6.2, this potential function strictly decreases at each cutting operation. Moreover, we have $\varphi(\mathscr{S})=2g-k$, and by the previous paragraph $\varphi(\mathscr{S}^{\prime})$ is at least $-(k+g+b)$. This proves (1).
@@
This section is devoted to the proof of Theorem 1.3. The proof technique is based on an irrelevant vertex method; we borrow ingredients to Kociumaka and Pilipczuk [22, Section 5] and to graph minor algorithms, but some new arguments are needed, in particular in the beginning of the proof of Proposition 8.1, and because we need to take singular vertices into account.
@@ 8.1 Finding a large planar part
Let $G$ be a graph with $n$ vertices and edges and $g\geq 2$ be an integer. In $2^{\text{poly}(g)}n$ time, we can do one of the following:
@@
We will use the following lemma, whose proof relies on some advanced results related to treewidth and tree decompositions.
@@
Let $G$ be a graph with $n$ vertices and edges, and $k\geq 2$. Then, one can compute either a rooted branch decomposition of $G$ of width polynomial in $k$ or a $(k\times k)$-grid minor of $G$, in time $2^{\text{poly}(k)}\cdot n$.
In the proof below, the polynomial in $k$ is shown to be $O(k^{9+\varepsilon})$, for any $\varepsilon>0$.
@@
In the same way as the branchwidth of a graph $H$ is the minimum width of a branch decomposition of $H$, the treewidth of $H$ is the minimum width of a tree decomposition of $H$. We do not define tree decompositions here, but it suffices to know that one can, in linear time, convert a branch decomposition of $G$ of width $w$ into a tree decomposition of $G$ of width at most $w+1$; and similarly in linear time we can convert a tree decomposition of $G$ of width $w$ into a branch decomposition of $G$ with width at most $13w/2-$ 1 [33, Theorem 5.1].
@@
By a result by Chekuri and Chuzhoy [9], we know that for some $d$ large enough, if $G$ has treewidth at least $k^{d}$, then it has a $(k\times k)$-grid minor. One can take $d=9+\varepsilon$, by a result by Chuzhoy and Tan [10].
@@
We will use an algorithm to approximate the treewidth by Bodlaender, Drange, Dregi, Fomin, Lokshtanov, Ma. Pilipczuk [5]: Given a graph $G$ with $n$ vertices and edges, and an integer $t$, in $O(2^{\text{poly}(t)}n)$ time, we either correctly report that the treewidth of $G$ is at least $t$, or compute a tree decomposition of width at most $6t$. In a first step, we apply it with $t=k^{d}$. If we get a tree decomposition of width at most $6k^{d}$, we immediately obtain, and return, a (rooted) branch decomposition of width at most $9k^{d}$, as desired. The rest of the proof focuses on the former case, in which the treewidth of $G$ is at least $k^{d}$.
@@
We will also use a result by Perkovic and Reed [32] (see also Sau, Stamoulis, and Thilikos [36, Proposition 28]), which implies the following: Given an input graph $H$ with $n$ vertices, we can compute in $O(n)$ time a minor $H^{\prime}$ of $H$ with $|V(H^{\prime})|\leq\frac{15}{16}|V(H)|$ such that the treewidth of $H^{\prime}$ is at least half the treewidth of $H$. (This is because contracting a matching of a graph divides its treewidth by at most two.)
@@
We alternately apply the algorithm by Bodlaender et al. with $t=2k^{d}$, and the algorithm by Perkovic and Reed, until Bodlaender’s algorithm computes a tree decomposition of width at most $12k^{d}$. This must happen at some point because the algorithm of Perkovic and Reed decreases the number of vertices geometrically. The step before this happens, the treewidth was at least $2k^{d}$, so at the end it is also at least $k^{d}$. We thus computed, in $O(2^{\text{poly}(k)}n)$ time, a minor $G^{\prime}$ of $G$ whose treewidth is between $k^{d}$ and $12k^{d}$, together with a tree decomposition of $G^{\prime}$ of width at most $12k^{d}$.
@@
By the choice of $d$, the graph $G^{\prime}$ has a $(k\times k)$-grid minor, which we can compute in $O(2^{\text{poly}(k)}n)$ time because we have a tree decomposition of width at most $12k^{d}$, using an algorithm by Adler, Dorn, Fomin, Sau, and Thilikos [1, Theorem 1] (the theorem is stated as a decision problem, but in Section 5 of the same paper it is explained how to explicitly find a minor). ∎
@@
Alternatively, with fewer and more standard tools, one could achieve this with a worse running time of $f(k)n^{2}$, for some computable function $f$. Indeed, we can first approximate the treewidth of $G$, e.g., by the algorithm by Bodlaender et al. [5]: Provided $d$ is large enough, in time $2^{\text{poly}(k)}n$, we either compute a tree decomposition of width at most $9dk^{d}$, and thus immediately obtain a (rooted) branch decomposition of width at most $6dk^{d}$ [33, Theorem 5.1], or correctly report that the treewidth of $G$ is at least $dk^{d}$, which by the result by Chekuri and Chuzhoy [9] implies the existence of a $(k\times k)$-grid minor, which we can compute in $f(k)n^{2}$ time, for some computable function $f$ [33, Algorithm 4.4].
@@
We first compute, in $O(g^{4}n)$ time, disjoint non-adjacent $(50g\lceil\sqrt{g}\rceil\times 50g\lceil\sqrt{g}\rceil)$-walls $W_{1},\dots,W_{g}$ of $W$, and the corresponding subdivisions $\dot{W}_{1},\dots,\dot{W}_{g}$ that are subgraphs of $\dot{W}$, in such a way that $W-W_{i}$ is connected for each $i$. For each $i$, we consider the subgraph $G_{i}$ of $G$ induced by the vertices $v$ of $G$ such that: (1) there exists a path from $v$ to $\dot{W}_{i}$; (2) every path from $v$ to $\dot{W}_{j}$, for some $j\neq i$, uses at least one vertex from $\dot{W}_{i}$. The graphs $G_{i}$ are pairwise disjoint, and we can compute each of them easily in $O(n)$ time using suitable traversals of $G$. We test the planarity of each of them in linear time [19]. If all of them are non-planar, we correctly report that $G$ has genus at least $g$ (Case 2). So without loss of generality, one of these graphs, say $G_{1}$, is planar, and our algorithm computes it.
@@
By 3-connectivity, $W_{1}$, and thus also $\dot{W}_{1}$, has a unique combinatorial embedding in the plane, up to symmetry and up to the choice of the outer (infinite) face; we consider the natural embedding of $W_{1}$ in which the outer face has the largest degree. An inner vertex of $W_{1}$ is one that is at distance at least 6 from the outer face in the natural embedding of $W_{1}$. Remark that each vertex of $W_{1}$ is a vertex of $\dot{W}_{1}$. We say that a vertex $u$ of $W_{1}$ is connected to the outside if there is, in $G$, a path whose vertices are, in this order, $u$, possibly some vertices of $\dot{W}_{1}$ but not of $W_{1}$, possibly some vertices of $G-\dot{W}_{1}$, and finally one vertex in $\dot{W}-\dot{W}_{1}$.

### v3

@@ 3.1 Some preprocessing
It is known that every graph can be embedded into a 3-book [15, Proposition 3.1]. So we can without loss of generality assume that $\mathscr{C}$ contains no 3-book. We remove all the isolated vertices of $\mathscr{C}$, and remove the same number of isolated vertices of $G$ (to the extent possible); this does not affect whether $G$ embeds into $\mathscr{C}$. We then replace each isolated vertex of $G$ with an isolated edge; since $\mathscr{C}$ has no more isolated vertices, this does not affect embeddability of $G$ into $\mathscr{C}$. Finally, for the same reason, if $G$ contains at least two connected components that are paths, we replace all these components with a single edge. ∎
@@ 3.2 Structure of 2-complexes without 3-book or isolated vertex
Let $\mathscr{C}$ be a 2-complex without 3-book or isolated vertex, and let $p$ be a vertex of $\mathscr{C}$. Following [15, Section 2.2], we describe the possible neighborhoods of $p$ in $\mathscr{C}$. A cone at $p$ is a cyclic sequence of triangles $t_{1},\ldots,t_{k},t_{1}$ ($k\geq 3$), all incident to $p$, such that, for each $i=1,\ldots,k$, the triangles $t_{i}$ and $t_{i+1}$ (where $t_{k+1}=t_{1}$) share an edge incident with $p$, and any other pair of these triangles have only $p$ in common. A corner at $p$ is an inclusionwise maximal sequence of distinct triangles $t_{1},\ldots,t_{k}$, all incident to $p$, such that, for each $i=1,\ldots,k-1$, the triangles $t_{i}$ and $t_{i+1}$ share an edge incident with $p$, any other pair of these triangles have only $p$ in common. An isolated edge at $p$ is an edge incident to $p$ but not incident to any triangle. The cones, corners, and isolated edges at $p$ form the link components at $p$.
@@ 3.3 Topological data structure for 2-complexes
Given a 2-complex $\mathscr{C}$ without 3-book or isolated vertex, described by vertices, edges, and triangles and their incidence relations, we can easily compute a representation of $\mathscr{C}$ in that data structure, in polynomial time in the number of simplices: Indeed, by ignoring the incidences created by vertices, we easily build a triangulation of the surface $\mathscr{S}$ (possibly disconnected, possibly with boundary) and a list of segments $S$; we then compute the topology of $\mathscr{S}$; finally, we mark the singular vertices, which are the vertices with several occurrences on $\mathscr{S}$ and/or on $S$. We remark that the size of the resulting data structure is at most linear in the number of simplices of the 2-complex $\mathscr{C}$, because, by Euler’s formula, any triangulated surface (possibly with boundary) with $k$ simplices has genus $O(k)$ and a number of boundary components that is $O(k)$. Thus, in the rest of this article, without loss of generality, we implicitly assume that $\mathscr{C}$ is given in the form of the above topological data structure. (Conversely, it is not hard to see that every 2-complex is homeomorphic to a 2-complex whose number of simplices that is linear in its size, but we will not need this fact.)
@@ 3.4 Proper and cellular graph embeddings on 2-complexes
Traditional data structures for graphs on surfaces handle graphs embedded cellularly; rotation systems [36] constitute one example of such a data structure. In order to have efficient algorithms, refined data structures, e.g., with graph-encoded maps [30] (see also [19, Section 2]), are needed. The basic element in the graph-encoded map is the flag, an incidence between a vertex, an edge, and a face of the graph. Three involutions allow to move from each flag to a nearby flag. Each flag contains a pointer to the underlying vertex, edge, and face.
One can easily extend such data structures to possibly non-cellular embeddings on surfaces [13, Section 2.2]. In this framework, one must store the topology of each face, which is not necessarily homeomorphic to a disk. Also, a face may have several boundary components; two-way pointers connect each face to one flag of each boundary component (or to an isolated vertex of the graph, if that boundary component is a single vertex); if a face is orientable and has several boundary components, then these pointers must induce a consistent orientation of these boundary cycles. It is important to remark that this data structure also allows to recover the topology of the underlying surface.
@@ 4 Partitioning graphs
  • Each face of $\Pi$ is labelled by an integer in $\{0,\ldots,k\}$ as follows: faces of $\Pi$ containing an edge in $E_{i}$ are labelled $i$, and the other ones are labelled 0. Since $\Pi$ contains the singular vertices of $\mathscr{C}$ and $\text{mid}(E_{1},\ldots,E_{k})$, and by the construction of the edges of $\Pi$, this labelling is well defined. We remark that this definition is valid also for faces of $\Pi$ that are parts of an isolated edge of $\mathscr{C}$.
@@
  • similarly, for each internal node $\nu$ of $\mathcal{B}$, the graph $\Pi(\Gamma,\nu)$ has at most $3(75c+27w)/2$ edges.
@@ 4.1 Monogons and bigons
A monogon of a graph $\Pi$ embedded on a 2-complex $\mathscr{C}$ is a face of $\Pi$ that is an open disk whose boundary is a single edge of $\Pi$ (a loop). Similarly, a bigon of $\Pi$ is a face of $\Pi$ that is an open disk whose boundary is the concatenation of two edges of $\Pi$ (possibly the same edge appearing twice). The following general lemma on graphs embedded on surfaces without monogons or bigons will be used below, and also in Section 8; some particular cases have been used before [9, Lemma 2.1].
@@
It is clear that the statement of the lemma holds whenever we are in one of the four above cases. So we can assume without loss of generality that each face of $\Pi^{\prime}$ is a triangle. Since $V(\Pi)=V(\Pi^{\prime})$ and $|E(\Pi)|\leq|E(\Pi^{\prime})|$, it suffices to prove the result for $\Pi^{\prime}$ instead of $\Pi$. Double-counting the incidences between edges and faces implies that the number of triangles $\tau$ satisfies $3\tau=2|E(\Pi^{\prime})|$; plugging this into Euler’s formula implies that $|V(\Pi^{\prime})|-|E(\Pi^{\prime})|/3=2-g$, so $|E(\Pi^{\prime})|=3g+3|V(\Pi^{\prime})|-6$, as desired. ∎
@@ 4.3 Rearranging $\Gamma$ with respect to an edge partition
Now, let us assume that $\Pi$ contains a sequence of bigons $B_{1},\ldots,B_{8}$ such that $B_{i}$ and $B_{i+1}$ share an edge for each $i$. So without loss of generality, we can assume that $B_{1}$ and $B_{5}$ are labelled $1$, $B_{3}$ and $B_{7}$ are labelled $2$, and the other bigons are labelled 0. We modify $\Gamma$ by exchanging the parts of $\Gamma$ inside $B_{3}$ and $B_{5}$; see Figure 4.5. The resulting embedding $\Gamma^{\prime}$ is also proper and cellular. This operation simplifies the endpoints $u$ and $v$ of these bigons. (If $u=v$, this is Case 2 from the definition of a simplification.)
@@
Finally, there remains to prove that, for each internal node $\nu$ of $\mathcal{B}$, the graph $\Pi(\Gamma,\nu)$ has at most $3(75c+27w)/2$ edges. Let $(E_{1},E_{2},E_{3})$ be the edge partition associated with $\nu$. By the claim we have just proved, it suffices to prove that the number of edges of this graph is at most half the sum of the numbers of edges of $\Pi(\Gamma,E_{1},E_{2}\cup E_{3})$, $\Pi(\Gamma,E_{1}\cup E_{2},E_{3})$, and $\Pi(\Gamma,E_{1}\cup E_{3},E_{2})$.
@@ 5.1 Bounding graphs
We compute our exhaustive list $\mathcal{L}_{\alpha}$ of sparse bounding graphs for $G_{\alpha}$ as follows. Initially, let this list be empty. Using Lemma 3.3, we enumerate all combinatorial maps $\Pi$ of graphs with at most $c+3w$ vertices and $3(75c+27w)/2$ edges properly embedded on $\mathscr{C}$ (possibly non-cellularly), with faces labelled 0, 1, 2, or 3, and such that the labels appearing on the vertices are exactly the vertices of the middle set of $\alpha$, $\beta$, or $\gamma$ (and each label appears exactly once). This takes $(c+w)^{O(c+w)}$ time. Whenever $\Pi_{1,3}\in\mathcal{L}_{\beta}$ and $(\Pi_{1,2})^{-}\in\mathcal{L}_{\gamma}$, we add $\Pi_{2,3}$ to $\mathcal{L}_{\alpha}$.
@@ 6 Reduction to proper cellular embeddings
The following result is not hard and essentially folklore (a related but slightly weaker result is provided by Matoušek, Sedgwick, Tancer, and Wagner [33, Lemma 3.1]), but we could not find a precise reference.
@@
    Case where $\gamma$ is non-separating but two-sided. This is only possible if $g\geq 2$. The cutting operation on $\mathscr{S}$ results in a single surface $\mathscr{S}^{\prime}$ with genus g2g-2. If $\mathscr{S}$ is orientable, then so is $\mathscr{S}^{\prime}$; otherwise, $\mathscr{S}^{\prime}$ is either orientable or non-orientable (unless of course $g=2$, in which case it is necessarily orientable, or $g$ is odd, in which case it is necessarily non-orientable). All surfaces $\mathscr{S}^{\prime}$ satisfying these constraints can be obtained.
@@
    Case where $\gamma$ is one-sided. This is only possible if $\mathscr{S}$ is non-orientable and $g\geq 1$. The cutting operation on $\mathscr{S}$ results in a single surface $\mathscr{S}^{\prime}$ with genus g1g-1, orientable or not (unless of course $g=1$, in which case it is orientable, or $g$ is even, in which case it is non-orientable). All surfaces $\mathscr{S}^{\prime}$ satisfying these constraints can be obtained.∎
@@
Let $\varphi(\mathscr{S}^{\prime})$ be equal to twice the total genus of $\mathscr{S}^{\prime}$ minus its number of connected components. By Lemma 6.2, this potential function strictly decreases at each cutting operation. Moreover, we have $\varphi(\mathscr{S})=2g-k$, and by the previous paragraph $\varphi(\mathscr{S}^{\prime})$ is at least (k+g+b)-(k+g+b). This proves (1).
@@
This section is devoted to the proof of Theorem 1.3. The proof technique is based on an irrelevant vertex method; we borrow ingredients from Kociumaka and Pilipczuk [27, Section 5] and from graph minor algorithms, but some new arguments are needed, in particular in the beginning of the proof of Proposition 8.1, and because we need to take singular vertices into account.
@@ 8.1 Finding a large planar part
Let $G$ be a graph with $n$ vertices and edges and $g\geq 2$ be an integer. In $O(2^{\text{poly}(g)}n)$ time, we can do one of the following:
@@
(Our proof implies a degree of $18+\varepsilon$, for any $\varepsilon>0$, for the polynomials in $g$ in the above statement.)
We will use the following lemma, whose proof relies on some advanced results related to treewidth and tree decompositions. It is probably folklore in some circles, and some similar ideas have been used recently [44, Section 3.2], but we could not find a reference.
@@
Let $G$ be a graph with $n$ vertices and edges, and $k\geq 2$. Then, one can compute either a rooted branch decomposition of $G$ of width polynomial in $k$ or a $(k\times k)$-grid minor of $G$, in time $O(2^{\text{poly}(k)}\cdot n)$.
(Our proof implies a degree of $9+\varepsilon$, for any $\varepsilon>0$, for the polynomials in $k$ in the above statement.)
@@
In the same way as the branchwidth of a graph $H$ is the minimum width of a branch decomposition of $H$, the treewidth of $H$ is the minimum width of a tree decomposition of $H$. We do not define tree decompositions here, but it suffices to know that one can, in linear time, convert a branch decomposition of $G$ of width $w$ into a tree decomposition of $G$ of width at most $w+1$; and similarly in linear time we can convert a tree decomposition of $G$ of width $w$ into a branch decomposition of $G$ of width at most 3w/213w/2-1 [40, Theorem 5.1].
@@
By a result by Chekuri and Chuzhoy [10], we know that for some large enough constants $c,d$, if $G$ has treewidth at least $ck^{d}$, then it has a $(k\times k)$-grid minor. One can take $d=9+\varepsilon$, by a result by Chuzhoy and Tan [11].
@@
We will use an algorithm to approximate the treewidth by Bodlaender, Drange, Dregi, Fomin, Lokshtanov, and Mi. Pilipczuk [6]: Given a graph $G$ with $n$ vertices and edges, and an integer $t$, in $2^{O(t)}n$ time, we either correctly report that the treewidth of $G$ is at least $t$, or compute a tree decomposition of width at most $10t$. In a first step, we apply it with $t=ck^{d}$. If we get a tree decomposition of width at most $10ck^{d}$, we immediately obtain, and return, a (rooted) branch decomposition of width at most $15ck^{d}$, as desired. The rest of the proof focuses on the former case, in which the treewidth of $G$ is at least $ck^{d}$.
@@
We will also use a result by Perkovic and Reed [39] (see also Sau, Stamoulis, and Thilikos [43, Proposition 28]), which implies the following: Given an input graph $H$ with $n$ vertices, we can compute in $O(n)$ time a minor $H^{\prime}$ of $H$ with $|V(H^{\prime})|\leq\frac{15}{16}|V(H)|$ such that the treewidth of $H^{\prime}$ is at least half the treewidth of $H$. (This is because contracting a matching of a graph divides its treewidth by at most two.)
@@
We alternately apply the algorithm by Bodlaender et al. with $t=2ck^{d}$, and the algorithm by Perkovic and Reed, until Bodlaender’s algorithm computes a tree decomposition of width at most $20ck^{d}$. This must happen at some point because the algorithm of Perkovic and Reed decreases the number of vertices geometrically. The step before this happens, the treewidth was at least $2ck^{d}$, so at the end it is also at least $ck^{d}$. We thus computed, in $O(2^{O(k)}n)$ time, a minor $G^{\prime}$ of $G$ whose treewidth is between $ck^{d}$ and $20ck^{d}$, together with a tree decomposition of $G^{\prime}$ of width at most $20ck^{d}$.
@@
By the choice of $c$ and $d$, the graph $G^{\prime}$ has a $(k\times k)$-grid minor, which we can compute in $O(2^{\text{poly}(k)}n)$ time because we have a tree decomposition of width at most $20ck^{d}$, using an algorithm by Adler, Dorn, Fomin, Sau, and Thilikos [1, Theorem 1] (the theorem is stated as a decision problem, but in Section 5 of the same paper it is explained how to explicitly find a minor). ∎
@@
Alternatively, with fewer and more standard tools, one could achieve this with a worse running time of $f(k)n^{2}$, for some computable function $f$. Indeed, we can first approximate the treewidth of $G$, e.g., by the algorithm by Bodlaender et al. [6]: Provided $c$ and $d$ are large enough, in time $2^{\text{poly}(k)}n$, we either compute a tree decomposition of width at most $10ck^{d}$, and thus immediately obtain a (rooted) branch decomposition of width at most $15ck^{d}$ [40, Theorem 5.1], or correctly report that the treewidth of $G$ is at least $ck^{d}$, which by the result by Chekuri and Chuzhoy [10] implies the existence of a $(k\times k)$-grid minor, which we can compute in $f(k)n^{2}$ time, for some computable function $f$ [40, Algorithm 4.4].
@@
We first compute, in $O(g^{4}n)$ time, disjoint non-adjacent $(50g\lceil\sqrt{g}\rceil\times 50g\lceil\sqrt{g}\rceil)$-walls $W_{1},\dots,W_{g}$ of $W$, and the corresponding subdivisions $\dot{W}_{1},\dots,\dot{W}_{g}$ that are subgraphs of $\dot{W}$, in such a way that $W-W_{i}$ is connected for each $i$. For each $i$, we consider the subgraph $G_{i}$ of $G$ induced by the vertices $v$ of $G$ such that: (1) there exists a path from $v$ to $\dot{W}_{i}$; (2) every path from $v$ to $\dot{W}_{j}$, for some $j\neq i$, uses at least one vertex from $\dot{W}_{i}$. The graphs $G_{i}$ are pairwise disjoint, and we can compute each of them easily in $O(n)$ time using suitable traversals of $G$. We test the planarity of each of them in linear time [24]. If all these disjoint graphs are non-planar, we correctly report that $G$ has genus at least $g$ (Case 2); this is easy to prove directly, and also follows from more general results, see, e.g., Miller [34, Theorem 1]. So without loss of generality, one of these graphs, say $G_{1}$, is planar, and our algorithm computes it.
@@
By 3-connectivity, $W_{1}$ has a unique combinatorial embedding in the plane, up to symmetry and up to the choice of the outer (infinite) face; we consider the natural embedding of $W_{1}$ in which the outer face has the largest degree. An inner vertex of $W_{1}$ is one that is at distance, in $W_{1}$, at least 6 from the vertices on the outer face in the natural embedding of $W_{1}$ (we emphasize that this is the distance measured in the graph $W_{1}$). Remark that each vertex of $W_{1}$ is a vertex of $\dot{W}_{1}$. We say that a vertex $u$ of $W_{1}$ is connected to the outside if there is, in $G$, a path whose vertices are, in this order, $u$, possibly some vertices of $\dot{W}_{1}$ but not of $W_{1}$, possibly some vertices of $G-\dot{W}_{1}$, and finally one vertex in $\dot{W}-\dot{W}_{1}$.